OpenStudy (grizzley95):

Find the distance between the points R(0, 5), S(12, 3) to the nearest tenth.

OpenStudy (anonymous):

x distance is 12 - 0 = 12 y distance is 3 - 5 = -2 Using Pythagorean's Theorem, 12^2 + (-2)^2 = c^2 144 + 4 = c^2 c^2 = 148 c = 12.1655 c = 12.2 where c is the length of a straight line connecting the two points.